OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

a = 1/4

OpenStudy (anonymous):

3a = 9 + 6 ->a = 15/3 = 5 => 4 * 5 -1 = 19

OpenStudy (anonymous):

3a = 15 a = 3 4(3) - 1 12 - 1 = 10

OpenStudy (anonymous):

a = 5 i mean 4(5) - 1 = 19

OpenStudy (anonymous):

your math is off shiroh